OpenStudy (anonymous):

Find the axis of symmetry of the graph of the function. f(x) = 2x^(2) - 12x + 19

OpenStudy (raden):

use the formula : x = -b/(2a)

OpenStudy (raden):

f(x) = 2x^(2) - 12x + 19 here known a = 2, b = -12, and c = 19

OpenStudy (raden):

x= -b/2a = -(-12)/(2*2) = 12/4 = 3

OpenStudy (anonymous):

thank you!!

OpenStudy (jhannybean):

your equation is \(\large y = 2x^2-12x+19\) Identify your a,b,c : \(\large a = 2 \ , \ b= -12 \ , \ c= +19\) formula for axis of symmetry: this is the point on your x-axis where the points on your parabola are symmetric. In order to find this, follow the formula, \(\large x= -\frac{b}{2a}\) We know our points, so just plug them into the formula \[\large x= -\frac{b}{2a}\]\[\large x= -\frac{-12}{2(2)} \]\[\large x= \frac{12}{4}\]\[\large x = 3\] your parabola on the left and right side of x=3 will be symmetrical.